Subject: minutes for APO 3.5m users committee meeting 10/21/13

From: Suzanne Hawley

Submitted: Fri, 25 Oct 2013 23:30:56 -0700 (PDT)

Message number: 1321 (previous: 1320, next: 1322 up: Index)

 		APO 3.5-m Users Committee Phonecon, 10/21/13

Attending:  Suzanne Hawley (and for Scott Anderson), Mark Klaene, Joanne 
Hughes Clark (Seattle Univ.), John Wisniewski (Univ. of Oklahoma), Jeff 
Bary (Colgate, for NAPG), Mike Crenshaw (Georgia State), Rene Walterbos, 
Jon Holtzman, John Bally, Mary Beth Kaiser, Dave Bowen (for Michael 
Strauss), Bill Ketzeback, Joe Huehnerhoff, and Bruce Gillespie

Absent:  Remy Indebetouw, Al Harper


Introduction of new Users Committee members:

Suzanne welcomed the newly appointed Users Committee members from several 
of our new leasing partners.  They are:

   o Jeff Bary - Colgate University, representing the Northeast Astronomy
                 Participation Group (NAPG members are Colgate, Bucknell,
                 Haverford, Hobart and William Smith, Middlebury and

   o Mike Crenshaw - Georgia State University

   o Joanne Hughes Clark - Seattle University

   o John Wisniewski - University of Oklahoma


User feedback and comments from institutional representatives:

o Washington (Suzanne Hawley) - Suzanne had nothing new to report.

o Chicago (Al Harper) - no report.

o New Mexico State (Rene Walterbos) - Rene had nothing new to report.

o Princeton (Dave Bowen) - Dave had nothing new to report.

o Colorado (John Bally) - John had nothing new to report.

o Virginia (Remy Indebetouw) - no report.

o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had nothing new to report. 
Mark thanked Mary Beth for providing us with useful comments on user 

o Georgia State (Mike Crenshaw) - Mike reported that Doug Gies and some 
students came to the site to use the telescope and had a very positive 

o Seattle Univ. (Joanne Hughes Clark) - Joanne noted that SPIcam seems to 
be performing better.

o NAPG (Jeff Bary) - Jeff said that there are several users within the 
NAPG that are already certified users of the 3.5-m.


Discussion of telescope/instruments report:

Mark highlighted that the yearly statistic for unscheduled down time came 
in at a commendable 0.7%, consistent with the good performance of the 
telescope in recent years.  For other highlights, DIS is working well 
although there is an ongoing issue with the DCam slit moving.  Agile is 
having rotator motor current problems that limit its range to +/- 180 
degrees.  TripleSpec is working OK, and SPIcam is seeing intermittent 
dark-current episodes, possibly due to ion pump interactions.  The echelle 
eCam images were streaking, which was fixed by working on the shutter and 
control computer.  NIC-FPS performance is unchanged.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                9/5/13 through 10/15/13

0) Overview

After 2 weeks of heavy rain and clouds, the typical dry and clear fall 
weather set in and the last few weeks have been productive.  We hosted a 
visit by a TV documentary crew on comet ISON through a NMSU observer.

1) Telescope

Telescope has been working very well since shutdown.  Last year?s recorded 
loss time due to equipment was 0.7%, another great year.

2) Instruments

DIS has been operating normally.  The only current issue is the DCam slit 
mask that appears to be moving around; this is being investigated.  The 
contamination on the red camera appears to be back (not unexpected) and 
will be addressed at next available time.

Agile is operational, with a slight problem with the rotator.  Currently 
the rotator is being restricted to +/- 180 degrees.  The instrument will 
be powered off normally and powered on and checked out 1 to 2 days before 
scheduled observing.

TripleSpec had no operational problems this period.  A problem with 
temperature metrology has come up and is being worked on with UVa team. 
A user also brought up an issue with data reduction that is being looked 
at by the UVa team.

SPIcam variable dark current has been isolated to the current ion pump 
type and controller.  Presently it is running with a different controller 
and problem is much reduced.  A different type of ion pump is on order.

Echelle is operational with no problems except for ECam which had a stuck 
shutter that is now repaired.

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issue.  NFS less than 9 is operational.


New APO imager:

Suzanne reminded the committee that our workhorse optical imager, SPIcam, 
is now ~15-years old and very difficult to maintain.  We've been looking 
to upgrade to a new imaging instrument for the past four years.  After 
several external approaches failed to materialize, we have decided to 
build the SPIcam replacement with in-house resources.  The scientific and 
functional requirements have been developed, and an internal team is at 
work to bring the camera to fruition in the coming year.

Bill Ketzeback said that we are holding a preliminary design review at APO 
on 30/31 October, and there is a web page with most of the review 
documentation posted.  Interested users should contact their users 
committee representative to gain access to the website.  The review will 
include outside experts, and feedback from APO staff and ARC users is most 

Suzanne added that members of this Users Committee should review and 
comment on the science project requirements posted at the website.  John 
Bally asked and it was confirmed that the new camera will accommodate 
narrow-band filters. Rene asked about the AR coating on the CCD.  Bill 
said that the plan is to be optimized for both UV and red response.  If 
users have a compelling reason to go to a different coating, we will 
consider it.  Suzanne said that if the PDR is successful, we will start 
the procurement of the hardware; a CDR will be held hopefully in early 
2014 and the ambitious plan is to have the instrument ready for 
commissioning within the next year.


Quarter 1 scheduling:

There is a large proposal from an outside group to lease 60 half nights in 
the next year, with 40 of them in quarter 1 2014.  Suzanne urged the Users 
Committee members to discuss this with their institutional schedulers as 
it is likely that we are going to accept the arrangement.  Many of the 
existing partners will be trading time from Q1 into later quarters to 
compensate.  The new leasing partners will get the time that has been 
agreed to in the signed MOUs.


User documentation:

Mark reported that Jon Holtzman has suggested that we put our user 
documentation under some kind of version control, such as SVN.  Jon also 
thinks that it would be a good time to create a user wiki, blog, forum, or 
other such modern communication tool so that users could post and discuss 
issues about the telescope, instrument, data reduction, etc.  Bill thought 
that a wiki or forum would be a good place for user discussion on data 
reduction, since that aspect of observing is not supported at APO.  Rene 
thought that a wiki would be a good thing, especially for reduction and 
archiving discussions.  Jeff and Joanne both felt that these kinds of 
improvements to our user documentation and communications would be 
especially good for the smaller user institutions.  Suzanne suggested that 
the Users Committee members talk to their respective user communities, and 
see if we can find volunteers who are interested in helping us set this 


Miscellaneous: apo35-general mailing list

Suzanne asked that the committee members from the new partner institutions 
send her the names and account information for individuals at their 
institutions that wish to be on the apo35-general mailing list.  The lead 
scientists are already subscribed.



[open from previous meeting]:

===> ACTION:  Find users to assist with instrument documentation.

 	Status:  Open.  Thanks to Michael Strauss and Mary Beth, who have 
sent useful comments.  We are waiting for comments on the infrared 
documentation especially from UVa and Colorado.

===> ACTION:  By 7 October, each institution to send 1 or 2 powerpoint 
slides to Suzanne and Mark highlighting 3.5-m science from the past year.

 	Status:  Closed, and thanks to those who sent materials.

[new from this meeting]:

===> ACTION:  Users committee members to query their communities for 
interest in and voluteers to help with setting up a 3.5m user wiki.


Next meeting:  The next Users Committee phonecon will be on Monday 2 
December at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1321 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to